A partnership has the following account balances: Cash, $90,000; Other Assets, $640,000; Liabilities, $344,000; Nixon (50 percent of profits and losses), $180,000; Cleveland (30 percent), $130,000; Pierce (20 percent), $76,000. The company liquidates, and $18,000 becomes available to the partners. Who gets the $18,000? Determine how much of this amount should be distributed to each partner. (Do not round intermediate calculations.)
Answer #1

The current net capital is = $90000 + $640000 - $344,000 = $386,000.

The company liquidates and $18,000 becomes available to partners, then the maximum potential company loss would be $386,000 - $18,000 = $368,000.
This potential loss is distributed to parterns in ratio 5:3:2
Nixon Distribution = $368,000 x 50% = $184,000
Cleveland Distribution = $368,000 x 30% = $110,400
Pierce Distribution = $368,000 x 20% = $73,600

In balance after distribution of loss, Nixon is the only one shown to have a potential deficit, therefore it fall to Cleveland and Pierce to cover the potential loss. This potential loss of Nixon is going to distribute to Cleveland and Pierce in 3:2
Cleveland Distribution = $4,000 x 3/5 = $2400
Pierce Distribution = $4,000 x 2/5 = $1600

Distribution statement is as follows:-

Now cash will be distribution to Cleveland and Pierce as $17,200 and $800 respectively.
Answer #2

Calculation for Distribution of $18,000

Step 1: Determine Profit/Loss Sharing Ratios

The partners share profits/losses as:

  • Nixon: 50%

  • Cleveland: 30%

  • Pierce: 20%


Step 2: Calculate Each Partner’s Share of the $18,000

Since the partnership is liquidating, the remaining cash is distributed based on their profit-sharing ratios.

  • Nixon (50%):
    18,000×0.50=9,000

  • Cleveland (30%):
    18,000×0.30=5,400

  • Pierce (20%):
    18,000×0.20=3,600

Step 3: Verify Against Capital Balances

Before finalizing, we check if any partner’s distribution exceeds their capital balance:

  • Nixon: 180,000(safe,since9,000 < $180,000)

  • Cleveland: 130,000(safe,since5,400 < $130,000)

  • Pierce: 76,000(safe,since3,600 < $76,000)


Since no restrictions apply, the full amounts are distributed.



Final Distribution

  • Nixon receives: $9,000

  • Cleveland receives: $5,400

  • Pierce receives: $3,600


answered by: anonymous